# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v4

@@ 2.3.2 Physical Property-embedded Machine Learning
Physical properties frequently encompass symmetries and conservation laws inherent in physical systems. Constructing a neural network that inherently respects a specific physical property enhances the likelihood of conserving relevant quantities, thereby improving the model’s predictions in terms of physical accuracy.
Conservation. Many real-world systems adhere to conservation laws related to mass, energy, momentum, or particle number, typically expressed through continuity equations. For example, in hydrology, it is the amount of water [167]. However, standard deep learning methods, such as CNNs and LSTMs, encounter challenges in maintaining conservation across layers or time steps. To address this issue, the mass-conserving LSTM (MC-LSTM) [168] extends the inductive bias of the LSTM to uphold these conservation laws, ensuring that mass input is conserved through modifications to the recurrent structure of the traditional LSTM. The main concept is to utilize memory cells from LSTMs as mass accumulators or storage units. In addition, an antisymmetrical continuous convolutional layer [169] in a hierarchical network is presented by enforcing the conservation of momentum with a hard constraint. The parameterization of deep neural networks [170] is tailored to meet the continuity equation, ensuring the adherence to fundamental conservation laws by creating divergence-free NNs. An implicit neural network layer [171] that incorporates the conservation of mass is proposed by using implicit differentiation. Furthermore, a structured approach to enforce nonlinear analytic constraints such as energy and mass conservation within NNs [172] is proposed. The concept of “conversion layers” is introduced within the architecture to transform nonlinearly constrained mappings into linearly constrained ones within NNs, achieving this without significantly compromising performance.
Symmetry has long been implicitly utilized in DL for designing networks with known invariances and equivariances. Let $f:X\rightarrow Y$ be a function and $G$ be a group. Assume $G$ acts on $X$ and $Y$. If $f(gx)=gf(x)$ for all $x\in X$ and $g\in G$, the function $f$ is $G$-equivariant. If $f(gx)=f(x)$ for all $x\in X$ and $g\in G$, the function $f$ is G-invariant.
Convolutional neural networks have revolutionized computer vision by utilizing translation equivariance. Similarly, symmetries are enforced in recurrent neural networks, graph neural networks, and capsule networks [12]. There exists a profound connection between symmetry and physics. For instance, Noether’s theorem establishes a relationship between conserved quantities and symmetry groups [173]. Moreover, symmetries in fluid dynamics [174] are employed in the design of equivariant networks. The Navier-Stokes equations and the heat equation exhibit invariance under the following transformations:
@@
Recently, incorporating symmetries of physics into NNs has been developed in the context of CNNs [175], GNNs, and neural operators [176].
First, an end-to-end ML-based deep potential–smooth edition (DeepPot-SE) [177] is developed, which is extensive, continuously differentiable, and scales linearly with system size while preserving all natural symmetries. The symmetry-preserving functions maintain translational, rotational, and permutational symmetries.
Second, there have been a number of GNN-based works, called geometrically equivariant graph neural networks [178], which leverage symmetry as an inductive bias in learning simulations. These models ensure that their outputs rotate, translate, or reflect in the same manner as their inputs, preserving symmetry.For example, EGNN [179] models interactions using invariant distances, calculated via inner products of relative positions, while GMN [180] extends to a multi-channel version with a stack of vectors. Subequivariant GNN [181] relaxes the condition of full equivariance to subequivariance, accommodating external fields like gravity while retaining the theoretical ability for universal approximation.
Thirdly, an integral neural operator architecture [176] is designed to automatically learn physical models with fundamental conservation laws guaranteed. This translation- and rotation-invariant neural operator replaces frame-dependent position information with invariant counterparts in kernel space, ensuring conservation laws of linear and angular momenta are upheld.
Causality. A fundamental pursuit in physics is to identify causal relationships. Integrating causality into ML promises to advance the comprehension of physical processes and bolster the robustness of ML models [182]. For example, causality-DeepONet [183] implements the physical causality in the DeepONet structure. Specifically, the causality-DeepONet is employed to learn operators that capture the response of buildings to earthquake ground acceleration. However, many challenges in causal embedding in ML remain unresolved. These include leveraging causality to enhance deep learning models, understanding system responses under interventions, disentangling complex and multiple processes, and designing environments to control physical phenomena. Additionally, a causal training algorithm [184] has been introduced to restore physical causality during PINNs model training by appropriately reweighting the PDE residual loss in each iteration of gradient descent. This straightforward adjustment enables PINNs to address complex problems such as the incompressible Navier-Stokes equations in turbulent regimes.
@@ 2.3.3 Physical Condition-embedded Machine Learning
There are many physical conditions in the physical system, such as boundary conditions and initial conditions in PDEs, boundaries of land use and land cover in the physical environment, and so on. Failing to satisfy these conditions can result in unstable models and non-physical solutions. Therefore, it is essential for ML models to adhere to these constraints to accurately capture underlying physical principles and provide reliable models for scientific and engineering applications. Specifically, for PDEs, three types of boundary conditions (BCs) that are widely used to model physical phenomena are (1) Dirichlet: fixed-value at the boundary; (2) Neumann: fixed-derivative at the boundary; and (3) periodic: equal values at the boundary. To integrate boundary conditions into NNs, a boundary enforcing operator network [185] is introduced. This network ensures the satisfaction of boundary conditions by modifying the operator kernel’s structure. The physics-embedded neural network [186] is developed by considering BCs in GNN.
@@ 2.4.1 Neural–numerical Hybrid Learning
Serial Hybrid has broad potential advantages for physical models and data-driven models. For example, NNs are used to handle data coming from irregular domains. A family of neural ODEs [187, 188] transforms traditionally discretized neuron layer depths into continuous equivalents by parameterizing the derivative of the hidden state with a neural network. The network’s output is then computed using a black-box differential equation solver, effectively merging NNs with numerical solvers for efficiency. In addition, the advantages of physics solvers (such as their scale-invariant properties) can be imparted into NNs. A novel hybrid training approach [189] that combines a scale-invariant physics solver with ML techniques is developed for solving inverse problems, like parameter estimation and optimal control. Gradients of NNs are replaced by the update computed from a higher-order solver that can encode the scale-invariant property of the physics.
@@
Parallel Hybrid methods are designed to develop fast, robust, and reliable solvers for solving PDEs. Physics-driven models are usually used as a supervised signal to train data-driven NNs. For example, a solver named HINTS (Hybrid, Iterative, Numerical, and Transferable Solver) [190] is proposed for differential equations. HINTS combines traditional relaxation techniques (such as Jacobi, Gauss-Seidel, conjugate gradient, multigrid methods, and their variants) with DeepONet. This approach offers faster solutions for a wide range of differential equations while maintaining machine-level accuracy. Additionally, a hybrid graph neural network [191] that merges a traditional graph convolutional network with an integrated differentiable fluid dynamics simulator has been developed. This hybrid model generalizes well to new scenarios and benefits from the significant speedup of neural network-based CFD predictions, significantly outperforming standalone coarse CFD simulations. A fast, differentiable hybrid approach [192] integrates a 2D direct numerical simulation for deformable solid structures with a physics-constrained neural network surrogate to capture fluid hydrodynamic effects.
@@
Complex Hybrid methods are flexible frameworks that can be used to address many problems in solving PDEs. For example, in order to address the scalability issue of solving PDEs on a large scale, a physics-aware downsampling method [193] is proposed by minimizing the distance between the solutions on the fine and coarse grids. Specifically, a serial hybrid method is designed to downsample fine grid terrain, by combining a numerical solver (such as finite difference) on coarse grid terrain and a downsampling neural network. Then, a parallel numerical solver on fine grid terrain is used to provide a supervised signal for the predictions on the coarse grids.
@@ 2.4.2 Neural–hydrology Hybrid Learning
The hydrological model is a standard representation of physical processes, serving as an input-output model used to simulate the evolution and dynamics of surface and groundwater storage, fluxes, and physical properties of the Earth [2]. Neural-hydrology hybrid learning integrates various hydrological models using machine learning methods to produce a final process-based prediction product. This approach has garnered increased attention recently due to its ability to enhance the interpretability of hydrological models, improve understanding of machine learning techniques, and leverage advances in computational resources and methods [30]. Here we list some recent neural–hydrology hybrid methods for hydrological variable prediction, such as streamflow and runoff.
@@
Parallel Hybrid. In a parallel hybrid architecture, data-driven models and dynamic models are integrated concurrently. This approach can employ a data-driven model to combine hydrological predictions in parallel. For instance, a hybrid approach combining conceptual and ML techniques has been proposed to enhance the accuracy of runoff simulations in snow-covered basins [198]. An end-to-end hybrid modeling approach is introduced to learn and predict the spatiotemporal variations of observed and unobserved hydrological variables globally [199]. This model integrates a dynamic neural network with a conceptual water balance model, constrained by water cycle observational products such as evapotranspiration, runoff, snow water equivalent, and changes in terrestrial water storage. The model accurately reproduces observed water cycle variations, and the relationships of runoff generation processes are well aligned with established understanding. Furthermore, a hybrid hydrological model [200] is used for global hydrological modeling, utilizing neural networks’ adaptability to represent uncertain processes within a framework grounded in physical principles, such as mass conservation. This hybrid model is simultaneously trained using a multi-task learning approach, indicating that it provides a novel data-driven perspective for modeling the global hydrological cycle and physical responses through machine-learned parameters. This method aligns with existing global modeling frameworks. Moreover, parallel hybrid methods have been shown to correct model biases [201].
@@
Complex Hybrid. In a complex structure, data-driven models can be integrated into the input space, output space, and interior of hydrological models in a complex hybrid method, according to different application requirements. Specifically, a general physics-AI approach [9] is proposed to improve AI geoscientific awareness, wherein temporal dynamic geoscientific models are included as a special process-wrapped recurrent neural network (denoted as P-RNN). In addition, the common NN layers that lack physics-based processes address those unrepresented by the P-RNN layer. Experiments with the physics-AI approach show that this complex hybrid model improves prediction accuracy, robustness in transferability, and the ability to infer unobserved processes.Additionally, a differentiable programming framework [202] is designed to parameterize, enhance, or replace the process-based model’s modules in a simple hydrologic model, HBV, for predicting hydrologic variables such as streamflow. Specifically, the LSTM outputs the physical parameters of the process-based model. Within the process-based model, certain components can be substituted with NNs and the structure is updatable. The framework is trained end-to-end, without intermediate supervisory data or labels. The primary loss function is computed between the output of the process-based model and the observed data.
2.5 Summary of Physics-aware Machine Learning
These four methods in the PaML community, including their corresponding benefits and drawbacks for scientific problems, are summarized in Table 4.
Table 4: Four Approaches to Physics-aware Machine Learning
Term Description Methods Benefits and drawbacks
Physical Data-guided Machine Learning (PDgML) PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data. Classical physical data-guided neural networks (spatio-temporal modeling, dynamic system modeling), deep operator networks, physics-discovery neural networks [Benefits] Only need simulations from physics model, fast inference, the discovery of the dynamic behavior;
Physics-informed Machine Learning (PiML) PiML is a widely used approach to incorporating physical constraints that can be trained from additional information obtained by enforcing the physical laws (for example, designing loss functions or regularization). Physics-informed neural networks (PINNs), physics-informed deep neural operators (PINOs) [Benefits] Incomplete models and imperfect data, strong generalization in the limited data, understanding deep learning by the optimization process, tackling high dimensionality; uncertainty quantification, data-physics-driven parameters discovery;
Physics-embedded Machine Learning (PeML) PeML is achieved by embedding physics in the model frameworks or modules. Physical equation-embedded machine learning (PEeML), physical property-embedded machine learning (PPeML), physical condition-embedded machine learning (PCeML) [Benefits] Interpretability of architecture and learning process, generalization;
Physics-aware Hybrid Learning (PaHL) PaHL directly combines pure physics-based models, such as numerical methods, climate, land, hydrology, and Earth system models with ML models. Depending on the hybrid approach, hybrid learning can be categorized as serial, parallel, or complex. Neural–numerical hybrid learning, neural–hydrology hybrid learning [Benefits] Can leverage both the flexibility of data-driven models and the interpretability and generalizability of physics-driven models, speed, and operational convenience;

### v5

@@ 2.3.2 Physical Property-embedded Machine Learning
Physical properties usually include conservation, symmetry, and causality of physical systems. By building a NN that inherently respects a given physical property, we thus make conservation of the associated quantity more likely and consequently the model’s prediction more physically accurate.
Conservation. Many real-world systems adhere to conservation laws related to mass, energy, momentum, or particle number, typically expressed through continuity equations. For example, in hydrology, it is the amount of water [165]. However, standard DL methods, such as CNNs and LSTMs, encounter challenges in maintaining conservation across layers or time steps. To address this issue, the mass-conserving LSTM (MC-LSTM) [166] extends the inductive bias of the LSTM to uphold these conservation laws, ensuring that mass input is conserved through modifications to the recurrent structure of the traditional LSTM. The main concept is to utilize memory cells from LSTMs as mass accumulators or storage units. In addition, an antisymmetrical continuous convolutional layer [167] in a hierarchical network is presented by enforcing the conservation of momentum with a hard constraint. The parameterization of deep neural networks [168] is tailored to meet the continuity equation, ensuring the adherence to fundamental conservation laws by creating divergence-free NNs. An implicit neural network layer [169] that incorporates the conservation of mass is proposed by using implicit differentiation. Furthermore, a structured approach to enforce nonlinear analytic constraints such as energy and mass conservation within NNs [170] is proposed. The concept of “conversion layers” is introduced in the architecture that transforms nonlinearly constrained mappings into linearly-constrained mappings within NNs without overly degrading performance.
Symmetry is implicitly leveraged in DL frameworks to design networks with invariant and equivariant properties. Let $f:X\rightarrow Y$ be a function and $G$ be a group. Assume $G$ acts on $X$ and $Y$. If $f(gx)=gf(x)$ for all $x\in X$ and $g\in G$, the function $f$ is $G$-equivariant. If $f(gx)=f(x)$ for all $x\in X$ and $g\in G$, the function $f$ is G-invariant.
CNNs enabled breakthroughs in computer vision by leveraging translation equivariance. Similarly, RNNs, GNN, and capsule networks all impose symmetries [12]. There is a deep connection between symmetries and physics. For instance, Noether’s theorem [171] establishes a relationship between conserved quantities and symmetry groups. Moreover, symmetries in fluid dynamics [172] are employed in the design of equivariant networks. The Navier-Stokes equations and the heat equation exhibit invariance under the following transformations:
@@
Recently, incorporating symmetries of physics into NNs has been developed in the context of CNNs [173], GNNs [174], and neural operators [175].
First, an end-to-end ML-based deep potential–smooth edition (DeepPot-SE) [176] is developed, which is extensive, continuously differentiable, and scales linearly with system size while preserving all natural symmetries. The symmetry preserving functions maintain translational, rotational, and permutational symmetries.
Second, there have been a number of GNN-based works, called geometrically equivariant graph neural networks [174], which leverage symmetry as an inductive bias in learning simulations. These models ensure that their outputs rotate, translate, or reflect in the same manner as their inputs, preserving symmetry.For example, EGNN [177] models interactions using invariant distances, calculated via inner products of relative positions, while GMN [178] extends to a multi-channel version with a stack of vectors. Subequivariant GNN [179] relaxes the condition of full equivariance to subequivariance, by considering external fields like gravity and retaining the theoretical ability for universal approximation.
Third, an integral neural operator architecture [175] is designed to learn physical models with fundamental conservation laws automatically guaranteed. The translation- and rotation-invariant neural operator is developed by replacing the frame-dependent position information with its invariant counterpart in the kernel space; consequently it abides by the conservation laws of linear and angular momenta. A group equivariant FNO [180] is introduced, featuring Fourier layers designed to encode symmetries by exhibiting equivariance to rotations, translations, and reflections within the neural operator architecture.
Causality. A fundamental pursuit in physics is to identify causal relationships. Integrating causality into ML promises to advance the comprehension of physical processes and bolster the robustness of ML models [181]. For example, causality-DeepONet [182] implements the physical causality in the DeepONet structure. Specifically, the causality-DeepONet is employed to learn operators that capture the response of buildings to earthquake ground acceleration. However, many challenges in causal embedding in ML remain unresolved. These include leveraging causality to enhance DL models, understanding system responses under interventions, disentangling complex and multiple processes, and designing environments to control physical phenomena. Additionally, a causal training algorithm [183] has been introduced to restore physical causality during PINNs model training by appropriately reweighting the PDE residual loss in each iteration of gradient descent. This straightforward adjustment enables PINNs to address complex problems such as the incompressible Navier-Stokes equations in turbulent regimes.
@@ 2.3.3 Physical Condition-embedded Machine Learning
There are many physical conditions in the physical system, such as boundary conditions and initial conditions in PDEs, boundaries of land use and land cover in the physical environment, and so on. Failing to satisfy these conditions can result in unstable models and non-physical solutions. Therefore, it is essential for ML models to adhere to these constraints. Physical condition-embedded ML establishes a connection with physical learning by embedding these physical conditions into ML frameworks or modules, ultimately accurately capturing underlying physical principles and providing reliable models for scientific and engineering applications. Specifically, for PDEs, three types of boundary conditions (BCs) that are widely used to model physical phenomena are (1) Dirichlet: fixed-value at the boundary; (2) Neumann: fixed-derivative at the boundary; and (3) periodic: equal values at the boundary. To integrate boundary conditions into NNs, a boundary enforcing operator network [184] is introduced. This network ensures the satisfaction of boundary conditions by modifying the operator kernel’s structure. The physics-embedded neural network [185] is developed by considering BCs in GNN.
@@ 2.4.1 Neural–numerical Hybrid Learning
Serial Hybrid has broad potential advantages for physical models and data-driven models. For example, NNs are used to handle data coming from irregular domains. A family of neural ODEs [186, 187] transforms traditionally discretized neuron layer depths into continuous equivalents by parameterizing the derivative of the hidden state with a NN. The network’s output is then computed using a black-box differential equation solver, effectively merging NNs with numerical solvers in a serial hybrid manner. In addition, the advantages of physics solvers (such as their scale-invariant properties) can be imparted into NNs. A novel hybrid training approach [188] that combines a scale-invariant physics solver with ML techniques is developed for solving inverse problems, like parameter estimation and optimal control. Gradients of NNs are replaced by the update computed from a higher-order solver that can encode the scale-invariant property of the physics. Furthermore, a serial hybrid unit combining a discrete physics solver and a NN [189] is proposed to solve complex fluid-structure interaction problems. The physics solver encodes the control PDEs in discrete form as a non-trainable component, which is seamlessly integrated with a trainable LSTM network to form the entire recurrent unit. The hybrid model is trained by considering both fluid and solid physics dynamics as loss functions. Experimental results demonstrate that the serial hybrid model outperforms purely data-driven neural models.
@@
Parallel Hybrid methods are designed to develop fast, robust, and reliable solvers for solving PDEs. Physics-driven models are usually used as a supervised signal to train data-driven NNs. For example, a solver named HINTS (Hybrid, Iterative, Numerical, and Transferable Solver) [190] is proposed for differential equations. HINTS combines traditional relaxation techniques (such as Jacobi, Gauss-Seidel, conjugate gradient, multigrid methods, and their variants) with DeepONet. This approach offers faster solutions for a wide range of differential equations while maintaining machine-level accuracy. A fast, differentiable hybrid approach [191] integrates a 2D direct numerical simulation for deformable solid structures with a physics-constrained neural network surrogate to capture fluid hydrodynamic effects. A hybrid model enhanced by the finite element method [192] is proposed to create a high-performance surrogate model for forward and inverse PDE problems. This model hybridizes the NN and the finite element method in parallel to output the physical quantities used to calculate the loss function (residual vector). Additionally, parallel hybrid allows numerical solvers to achieve high accuracy on coarse grids, significantly reducing computational time. For instance, a deep neural network multigrid solver [193] enhances the finite element method on coarse grids by utilizing fine-scale information from NNs. This hybrid method is realized through the interaction between the finite element method and NNs on local grids.
@@
Complex Hybrid (Coupling Serial and Parallel Hybrids) methods are flexible frameworks that can be used to address many problems in solving PDEs. As shown in Fig. 4(c) top, data-driven models can be integrated into the input space by serial hybrid and the internal space by parallel hybrid. For example, in order to address the scalability issue of solving PDEs on a large scale, a physics-aware downsampling method [194] is proposed by minimizing the distance between the solutions on the fine and coarse grids. Specifically, a serial hybrid method is designed to downsample fine grid terrain, by combining a numerical solver (such as finite difference) on coarse grid terrain and a downsampling neural network. Then, a parallel numerical solver on fine grid terrain is used to provide a supervised signal for the predictions on the coarse grids. In addition, as shown in Fig. 4(c) below, data-driven models can be integrated into the internal space by parallel hybrid and the output space by serial hybrid. For instance, a hybrid graph neural network [195] that merges a traditional graph convolutional network (GCN) with an integrated differentiable fluid dynamics simulator has been developed. Through parallel hybrid, fine grid parameters are input to the GCN to obtain the graph representation, while the coarse grid parameters are input to the fluid dynamics simulator to generate simulation results. The upsampled simulation results from the fluid dynamics simulator are subsequently fed into a GCN through serial hybrid, ultimately predicting the desired output values. This complex hybrid model generalizes well to new scenarios and benefits from the significant speedup of neural network-based CFD predictions, significantly outperforming standalone coarse CFD simulations.
@@ 2.4.2 Neural–hydrology Hybrid Learning
The hydrological model is a standard representation of physical processes, serving as an input-output model used to simulate the evolution and dynamics of surface and groundwater storage, fluxes, and physical properties of the Earth [2]. Neural-hydrology hybrid learning integrates various hydrological models using ML methods to produce a final process-based prediction product. This approach has garnered increased attention recently due to its ability to enhance the interpretability of hydrological models, improve understanding of ML techniques, and leverage advances in computational resources and methods [30]. Here we list some recent neural–hydrology hybrid methods for hydrological variable prediction, such as streamflow and runoff.
@@
Parallel Hybrid. In a parallel hybrid architecture, data-driven models and physics-driven models are integrated concurrently. This approach can employ a data-driven model to combine hydrological predictions in parallel. For instance, a hybrid approach combining conceptual and ML techniques has been proposed to enhance the accuracy of runoff simulations in snow-covered basins [200]. An end-to-end hybrid modeling approach is introduced to learn and predict the spatiotemporal variations of observed and unobserved hydrological variables globally [201]. This model integrates a dynamic neural network with a conceptual water balance model, constrained by water cycle observational products such as evapotranspiration, runoff, snow water equivalent, and changes in terrestrial water storage. The model accurately reproduces observed water cycle variations, and the relationships of runoff generation processes are well aligned with established understanding. Furthermore, a hybrid hydrological model [202] is used for global hydrological modeling, utilizing neural networks’ adaptability to represent uncertain processes within a framework grounded in physical principles, such as mass conservation. This hybrid model is simultaneously trained using a multi-task learning approach, indicating that it provides a novel data-driven perspective for modeling the global hydrological cycle and physical responses through machine-learned parameters. This method aligns with existing global modeling frameworks. Moreover, parallel hybrid methods have been shown to correct model biases [203].
@@
Complex Hybrid (Coupling Serial and Parallel Hybrids). In a complex structure, data-driven models can be integrated into the input space, output space, and interior of hydrological models in a complex hybrid method, according to different application requirements. First, as shown in Fig. 4(c) top, data-driven models can parameterize hydrological models in the input space through serial hybrid. Simultaneously, data-driven models can effectively represent the internal dynamics of hydrological processes by integrating with traditional hydrological models. For example, a general physics-AI approach [9] is proposed to improve AI geoscientific awareness, wherein temporal dynamic geoscientific models are included as a special process-wrapped recurrent neural network (denoted as P-RNN). Common NN layers that lack physics-based processes address those unrepresented by the P-RNN layer. This combination improves the representation of rainfall-runoff processes. Additionally, a parameterization pipeline using NN layers through serial hybrid maps region-dependent attributes (e.g., soil properties and topography) onto the P-RNN and common NN layers. Experiments with the physics-AI approach show that this complex hybrid model improves prediction accuracy, model transferability, and the ability to infer unobserved processes. Additionally, a differentiable programming framework [204] is designed to parameterize, enhance, or replace the process-based model’s modules in a simple hydrologic model, HBV, for predicting hydrologic variables such as streamflow. Specifically, the LSTM outputs the physical parameters of the process-based model via serial hybrid method. Within the process-based model, certain components can be substituted with NNs and the structure is updatable. The framework is trained end-to-end, without intermediate supervising data or labels. The primary loss function is computed between the output of the process-based model and the observed data. Second, as shown in Fig. 4(c) below, data-driven models not only effectively capture the internal dynamics of hydrological processes by integrating with traditional hydrological models but also enhance model prediction performance in the output space as a post-processing layer. For instance, a process-driven DL model [205] is developed to enhance the process awareness of DL models. A conceptual hydrological model is integrated into an RNN cell to represent runoff sub-processes effectively, and an LSTM is used as a post-processor layer to output relevant system responses. Experimental results demonstrate that this complex hybrid model better characterizes the rainfall-runoff relationship, improving prediction performance.